### 13 Mar 2026 v1

@@ Abstract
Active learning (AL) has emerged as a powerful paradigm for accelerating materials discovery by iteratively steering experiments toward the most promising candidates, reducing costly synthesis-and-characterization cycles. However, current AL relies predominantly on Gaussian Process (GP) and Random Forest (RF) surrogates with complementary limitations: GP underfits complex composition–property landscapes due to rigid kernel assumptions, while RF produces unreliable uncertainty estimates in small-data regimes, precisely where most materials datasets (500 samples) reside. Here we propose In-Context Active Learning (ICAL), replacing conventional surrogates with TabPFN, a transformer-based foundation model pre-trained on millions of synthetic tasks to meta-learn a universal prior over tabular data. TabPFN performs principled Bayesian inference in a single forward pass without dataset-specific retraining, delivering well-calibrated predictive uncertainty where GP and RF fail most severely. Benchmarked against GP and RF across 10 materials datasets spanning copper alloy hardness and electrical conductivity, bulk metallic glass-forming ability, and crystal lattice thermal conductivity, TabPFN wins on 8 out of 10 datasets, achieving a mean saving of 52% in extra experiments/evaluations relative to GP and 29.77% relative to RF. Cross-validation analysis confirms that TabPFN’s advantage stems from superior uncertainty calibration, achieving the lowest Negative Log-Likelihood and Area Under the Sparsification Error curve among all surrogates. This demonstrates that a pre-trained foundation model can serve as a highly effective surrogate for active learning-based materials discovery.
Keywords deep learning one-class classifier materials discovery graph neural network deep learning
@@ 1 Introduction
The breadth of AL’s impact as a computation engine driving materials discovery is substantial and growing. In functional materials, Yuan et al. demonstrated accelerated discovery of large electrostrains in BaTiO3-based piezoelectrics [54], and Bassman et al. applied AL to layered materials design [6]. In electrocatalysis, Moon et al. used AL to discover a champion four-metal perovskite oxide for oxygen evolution [33], Suvarna et al. streamlined development of high-performance catalysts for higher alcohol synthesis [41], and Mok et al. identified electrocatalysts for CO2 reduction [15]. For alloy systems, Kim et al. searched for optimal multi-metallic alloy catalysts [21], and Rao et al. developed closed-loop AL frameworks for high-entropy Invar alloys [38]. Beyond metals, Kunkel et al. applied AL to organic semiconductor discovery [23]. AL has also transformed chemistry through self-driving laboratories (SDLs), which automate experimental planning and execution to accelerate the research pipeline in both chemistry and materials science [45, 9, 42, 1]. In drug design, Bayesian AL frameworks have been applied to drug screening [46] and regression-based AL has been used to accelerate large-scale virtual library docking [32]. Across all these domains, the common thread is that AL enables order-of-magnitude reductions in the number of experiments required relative to conventional design-of-experiments or random search [18, 29, 24, 31, 22].
@@
Motivation: TabPFN as a Foundation Model Surrogate for Active Learning
To address the dual limitations of existing surrogate models, we propose using TabPFN [17]— a foundation model for tabular data — as the surrogate in a Bayesian active learning framework for materials discovery. TabPFN is a transformer-based model that has been meta-trained via in-context learning on a large corpus of synthetic tabular datasets to approximate Bayesian inference over small tabular data problems . Unlike GP, RF, and NN surrogates that are fit from scratch on each new dataset, TabPFN leverages a pre-trained prior over functions and performs inference in a single forward pass without any gradient-based optimization on the target dataset. Critically, TabPFN outputs a full predictive distribution — not merely a point estimate with a heuristic variance proxy — because it was explicitly meta-trained to represent posterior uncertainty over tabular inputs. This makes it a principled probabilistic surrogate that is both expressive enough to capture complex composition–property relationships and capable of producing reliable uncertainty estimates in the small-data, high-dimensional settings that characterize experimental materials discovery.
@@ 3.1 Performance evaluation of ICAL in active-learning based materials discovery
Figure 2 presents the active learning performance on the cu_hardness dataset, where the target property is copper alloy hardness and the metric is the number of extra evaluations required to identify the global optimum beyond the initial training set (lower is better). TabPFN with UCB acquisition achieves a 90% win rate against both GP (UCB-norm) and RF (EI-norm) baselines — the highest win rate observed across all datasets tested — demonstrating a remarkably consistent and large performance advantage. As shown in panels (a) and (b), GP requires up to 110 extra evaluations at low init_ratios (0.05–0.10), while TabPFN reduces this to 38 and 5 respectively, representing a 3–20 reduction in experimental effort in the most data-scarce regime. This advantage persists monotonically as init_ratio increases, with TabPFN consistently requiring fewer than 20 extra evaluations across all conditions beyond init_ratio , while GP remains elevated at 25–50 evaluations through mid-range init_ratios. The RF comparison (panels c and d) reveals an even more problematic failure mode: RF (EI-norm) exhibits severe instability, with multiple sharp spikes reaching 85 extra evaluations at init_ratio and 65 at init_ratio . These spikes directly reflect RF’s well-known weakness in small-data regimes, where its heuristic ensemble variance produces unreliable uncertainty estimates that mislead the EI acquisition function into querying uninformative or redundant compositions. Critically, panel (d) shows that RF’s error bars are extraordinarily large — extending to 155 at init_ratio — meaning that in a single experimental campaign, a practitioner using RF could require over 150 additional synthesis-and-characterization cycles to find the optimal hardness alloy, at a potential cost exceeding $300,000 at conservative per-sample estimates. In stark contrast, TabPFN maintains both low mean extra evaluations and tight error bars throughout, confirming that its meta-trained in-context posterior provides not only superior average efficiency but also reliable, reproducible convergence that is essential for practical materials discovery workflows.
@@
Figure 3 (a) shows the active learning results on the cu_electric dataset targeting electrical conductivity, where TabPFN (EI) achieves a 75% win rate against both GP (EI) and RF (EI) baselines. Unlike the cu_hardness dataset where TabPFN’s advantage was consistent across all init_ratios, a notable exception emerges here at very low initial training set sizes (init_ratio ): TabPFN requires 80–83 extra evaluations, slightly exceeding GP’s 38–79 and RF’s 22–42 in this regime. This reversal suggests that TabPFN’s in-context posterior requires a minimum amount of task-specific data to form a reliable representation of the conductivity landscape before its modeling advantage over GP and RF materializes. Beyond init_ratio , however, TabPFN’s superiority becomes pronounced and sustained: GP remains persistently elevated at 40–85 extra evaluations with a notable peak at init_ratio (85), while TabPFN drops to below 15 evaluations across the same range. RF (panel c) exhibits its characteristic erratic behavior with a sharp spike to 105 at init_ratio , further confirming the unreliability of ensemble-variance-based uncertainty quantification in sparse data regimes. Both curves converge toward zero at init_ratio , as expected when the training set covers nearly the entire candidate space. The crossover behavior at low init_ratios is an important practical finding: it suggests that a minimum initial dataset size — approximately 15–20% of the candidate pool — should be collected before deploying TabPFN-based active learning for electrical conductivity optimization.
@@
Figure 3 (b) (d) presents results on the glass_DS3 dataset with magpie features targeting glass forming capability of alloys. TabPFN (UCB) achieves its highest win rate across all datasets tested: 95% against GP (EI) and 80% against RF (UCB-norm), suggesting that the glass forming capability landscape is particularly well-suited to TabPFN’s in-context Bayesian inference. Identifying glass-forming alloys is notoriously challenging due to the complex, nonlinear interplay between compositional features and the tendency to avoid crystallization — precisely the kind of structure that GP’s rigid kernel assumptions fail to capture. As shown in the upper panel, GP requires 60–90 extra evaluations across nearly the entire init_ratio range of 0.05–0.65, while TabPFN stays consistently below 45 throughout, representing a 2–4 reduction in the number of alloy synthesis and characterization experiments needed. This is practically significant given that producing and validating a glass-forming alloy — involving arc melting, melt spinning, and structural characterization via XRD to confirm amorphous phase formation — is among the more expensive experimental workflows in alloy design. The RF comparison (lower panel) tells a more nuanced story: RF and TabPFN perform comparably in the mid-range (init_ratio –), with curves frequently crossing, which accounts for RF’s 20% win rate — the highest any baseline achieves across all datasets. Nevertheless, TabPFN maintains a decisive advantage at low init_ratios where it matters most, with RF spiking to 77 extra evaluations at init_ratio while TabPFN requires only 37, and TabPFN’s overall smoother trajectory confirming more reliable convergence toward the optimal glass-forming composition.
@@
Figure 4 presents active learning results on the ltc_conc dataset targeting the lattice thermal conductivity (LTC) of crystal materials described by concentration features. Several distinctive and noteworthy patterns emerge that set this dataset apart from the others. TabPFN (UCB_hybrid) achieves a 75% win rate against GP (UCB) and a remarkable 95% win rate against RF (UCB) — the highest RF win rate across all datasets tested — indicating that RF is particularly ill-suited for navigating the LTC–composition landscape of crystal materials when only concentration features are used. The most striking observation in this dataset is the dramatic collapse to fewer than 10 extra evaluations beyond init_ratio : both GP and TabPFN achieve this near-perfect convergence all the way to init_ratio (panels a and b). This suggests that the LTC landscape of crystal materials parameterized by concentration features is relatively low-complexity and well-structured — once roughly 20% of the compositional space is sampled, the global optimum becomes straightforward to locate for both GP and Tabpfn surrogates. This is physically intuitive: LTC in crystals is governed primarily by phonon scattering mechanisms, which are strongly correlated with atomic mass contrast and bond stiffness — both of which vary smoothly with elemental concentration — producing a relatively smooth and navigable property landscape compared to mechanical properties or glass-forming ability. However, this result is still surprising as accurate prediction of LTC is extremely difficult by itself, showing the different ML model capability required for AL.
@@
The RF comparison (panels c and d) is even more revealing. RF (UCB) fails catastrophically and persistently: it requires 100–160 extra evaluations across the entire range from init_ratio to (panel d), never converging to fewer than 10 extra evaluations even when 50% of the candidate crystal compositions have been labeled. This is a qualitatively different failure mode from GP — RF does not merely need more data to converge, it fails to converge efficiently at all in this crystal composition feature space. The large error bars on RF in panel (d) further confirm that RF’s UCB-guided search is highly sensitive to random seed, producing wildly inconsistent outcomes across runs — a serious liability when synthesizing and characterizing crystal materials carries significant cost and effort. TabPFN, by contrast, converges to fewer than 10 extra evaluations by init_ratio with tight variance, demonstrating that the UCB_hybrid acquisition combined with TabPFN’s calibrated uncertainty is both efficient and reproducible for crystal LTC optimization.
@@
A particularly important finding emerges when comparing concentration-based versus Magpie feature sets for the copper and LTC datasets. For cu_electric, TabPFN with concentration features requires only 30.1 extra evaluations on average, compared to 58.5 (GP) and 37.5 (RF) — reductions of 49% and 20% respectively. However, switching to the high-dimensional Magpie feature set (cu_magpie_electric, 120 features) reverses the winner entirely: GP becomes best at 82.5 extra evaluations, while TabPFN degrades to 83.9 — nearly indistinguishable from GP and worse than both concentration-feature models. A similar pattern holds for cu_hardness: TabPFN achieves only 8.8 extra evaluations with concentration features (reductions of 72% vs. GP and 53% vs. RF), but with Magpie features (cu_magpie_hardness) GP wins at 21.9, and all three models perform substantially worse than their concentration-feature counterparts. The same trend is observed for LTC crystal materials: ltc_conc gives TabPFN a win at 39.9 extra evaluations (reductions of 29% vs. GP and 35% vs. RF), whereas ltc_magpie produces a dramatically inflated GP count of 243.88 — a 6 degradation relative to ltc_conc — before TabPFN partially recovers to 41.70. These results consistently indicate that high-dimensional Magpie features introduce severe overfitting in GP’s kernel estimation and destabilize RF’s ensemble variance, undermining the uncertainty quantification that drives acquisition function decisions. TabPFN is more robust to this feature dimensionality increase, but still performs best with the leaner concentration feature representation. This additionally shows that the AL needs different model capability than the regular regression performance to achieve high performance: the magpie feature can obtain better LTC prediction models, but not necessarily better AL performance.
@@
In contrast, the glass-forming alloy datasets reveal a strikingly different trend with respect to feature complexity. All three glass datasets are won by TabPFN, with mean extra evaluations of 48.8 (glass_ds1, concentration features), 32.6 (glass_ds2, physical features), and 28.2 (glass_ds3, Magpie features) respectively. Notably, performance improves monotonically as feature richness increases — the opposite of what is observed for copper and LTC datasets. The Magpie feature set yields the best result for glass-forming ability prediction, reducing extra evaluations by 58% relative to GP (67.5) and 11% relative to RF (31.6). This suggests that glass-forming ability is governed by a more complex, nonlinear composition–property relationship that genuinely benefits from richer physicochemical descriptors: predicting whether an alloy will form an amorphous phase depends on subtle electronic structure, atomic size mismatch, and thermodynamic factors that concentration features alone cannot adequately encode, but which Magpie descriptors partially capture. This stands in direct contrast to copper hardness and LTC, where the property landscape is smoother and high-dimensional Magpie features introduce more noise than signal. The feature-set sensitivity thus appears to be material-class dependent, and the consistent superiority of TabPFN across all three glass datasets — regardless of feature type — further demonstrates its robustness to feature space variations. Finally, ltc_structure is the only dataset where RF wins (84.1 extra evaluations for TabPFN vs. 55.05 for RF), suggesting that structural features present a unique inductive bias that RF’s decision-tree architecture exploits more effectively than TabPFN’s tabular meta-prior in this specific setting.
@@
For the lattice thermal conductivity (LTC) datasets, TabPFN with concentration features (ltc_conc) achieves savings of 29.4% vs. GP and 35.4% vs. RF and is selected as the best-performing feature set for the mean saving calculation, as the Magpie variant (ltc_magpie) offers only marginal additional improvement over RF (5.6%) despite a dramatic 82.9% reduction vs. GP — the latter being inflated by GP’s catastrophic failure with high-dimensional Magpie features rather than a genuine improvement in the active learning loop. The ltc_structure dataset is the sole case where TabPFN underperforms, with RF winning at 55.05 extra evaluations vs. TabPFN’s 84.1, suggesting that structural features encode an inductive bias that RF’s decision-tree architecture exploits more effectively than TabPFN’s tabular meta-prior; this dataset is excluded from the mean saving calculation accordingly. Taken together, the feature set analysis reveals a consistent principle: TabPFN performs best with the leanest feature representation that adequately captures the relevant physics of the target property, while high-dimensional feature sets only benefit TabPFN when the property landscape is sufficiently complex to warrant richer descriptors, as in the case of glass-forming ability.

### 17 Mar 2026 v2

@@ Abstract
Active learning has emerged as a powerful strategy for accelerating materials discovery by guiding expensive synthesis and characterization experiments toward the most informative candidates. However, its effectiveness critically depends on the surrogate model used to estimate both predictions and uncertainty. In experimental materials science, datasets are typically extremely small—often containing only tens to hundreds of samples—making reliable surrogate modeling challenging. Gaussian processes provide principled uncertainty estimates but often underfit complex composition–property relationships, while high-capacity models such as random forests and neural networks can capture nonlinear relationships but frequently produce unreliable uncertainty estimates in small-data regimes.
Here we introduce In-Context Active Learning (ICAL), a framework that replaces conventional surrogate models with a foundation model trained to approximate Bayesian inference over tabular data. By leveraging a meta-learned prior derived from millions of synthetic regression tasks, the model performs probabilistic inference in a single forward pass without dataset-specific retraining, enabling expressive prediction and calibrated uncertainty in small-data settings.
We evaluate ICAL across ten benchmark datasets spanning copper alloys, bulk metallic glasses, crystal lattice thermal conductivity, and electrolyte materials. ICAL consistently reduces the number of evaluations required to identify optimal materials, achieving a mean reduction of 52% relative to Gaussian-process-based active learning and 29.77% relative to random-forest baselines. Mechanistic analysis shows that these improvements arise from superior uncertainty calibration, allowing acquisition functions to more effectively balance exploration and exploitation during the discovery process.
These results demonstrate that foundation models can serve as effective surrogate models for data-efficient active learning, suggesting a new paradigm for accelerating discovery in small-data experimental sciences such as materials science, chemistry, biology, and other experimental disciplines characterized by small datasets and expensive evaluations.
Keywords active learning materials discovery foundation model small datasets
@@ 1 Introduction
Since the earliest active learning materials discovery studies in 2016 [54], the breadth of AL’s impact as a computation engine driving materials discovery is substantial and growing. In functional materials, Yuan et al. demonstrated accelerated discovery of large electrostrains in BaTiO3-based piezoelectrics [55], and Bassman et al. applied AL to layered materials design [6]. In electrocatalysis, Moon et al. used AL to discover a champion four-metal perovskite oxide for oxygen evolution [33], Suvarna et al. streamlined development of high-performance catalysts for higher alcohol synthesis [41], and Mok et al. identified electrocatalysts for CO2 reduction [15]. For alloy systems, Kim et al. searched for optimal multi-metallic alloy catalysts [21], and Rao et al. developed closed-loop AL frameworks for high-entropy Invar alloys [38]. Beyond metals, Kunkel et al. applied AL to organic semiconductor discovery [23]. AL has also transformed chemistry through self-driving laboratories (SDLs), which automate experimental planning and execution to accelerate the research pipeline in both chemistry and materials science [45, 9, 42, 1]. In drug design, Bayesian AL frameworks have been applied to drug screening [46] and regression-based AL has been used to accelerate large-scale virtual library docking [32]. Across all these domains, the common thread is that AL enables order-of-magnitude reductions in the number of experiments required relative to conventional design-of-experiments or random search [18, 29, 24, 31, 22].
@@
This tension between predictive capacity and uncertainty reliability represents a central challenge in active-learning-driven materials discovery. We hypothesize that foundation models trained to approximate Bayesian inference over tabular data can serve as universal surrogate models for small-data experimental discovery, resolving this long-standing trade-off and enabling more efficient experimental exploration.
Motivation: Foundation Model Surrogate for Active Learning
To address this limitation, we propose In-Context Active Learning (ICAL), a framework that replaces conventional surrogate models with a foundation model trained to approximate Bayesian inference over tabular data. Unlike traditional surrogates that must be trained from scratch for each dataset, the model leverages a meta-learned prior derived from large-scale synthetic training tasks, enabling expressive prediction and calibrated uncertainty even in extremely small-data regimes typical of experimental discovery.
Our ICAL algorithm uses TabPFN as the surrogate model, which is a transformer-based foundation model that has been meta-trained via in-context learning on a large corpus of synthetic tabular datasets to approximate Bayesian inference over small tabular data problems [17]. Unlike GP, RF, and NN surrogates that are fit from scratch on each new dataset, TabPFN leverages a pre-trained prior over functions and performs inference in a single forward pass without any gradient-based optimization on the target dataset. Critically, TabPFN outputs a full predictive distribution — not merely a point estimate with a heuristic variance proxy — because it was explicitly meta-trained to represent posterior uncertainty over tabular inputs. This makes it a principled probabilistic surrogate that is both expressive enough to capture complex composition–property relationships and capable of producing reliable uncertainty estimates in the small-data, high-dimensional settings that characterize experimental materials discovery.
@@
We hypothesize that foundation models trained to approximate Bayesian inference over tabular data can serve as universal surrogate models for small-data experimental discovery, resolving the long-standing trade-off between predictive capacity and reliable uncertainty estimation that limits conventional active learning approaches. This limitation arises from the small-data nature of experimental science. Foundation models offer a potential solution to this problem by providing meta-learned priors over functions.
@@ 2.3 Benchmark Datasets
We evaluate our active learning framework on ten molecular property prediction datasets derived from ChEMBL, spanning a diverse range of ADMET (absorption, distribution, metabolism, excretion, and toxicity) endpoints critical to drug discovery. The datasets vary substantially in size, ranging from 642 molecules (FreeSolv, hydration free energy) to 7,385 molecules (LD50, acute oral toxicity), and include: Caco-2 permeability (910), clearance microsome (1,102), clearance hepatocyte (1,213), VDSS volume of distribution (1,130), PPBR plasma protein binding rate (1,614), lipophilicity (4,200), LD50 (7,385), half-life (667), FreeSolv (642), and SARS-CoV-2 3CLpro inhibition (880). Each molecule is represented by 210 RDKit physicochemical descriptors, providing a unified feature space across all tasks. This collection covers a broad spectrum of biological and physicochemical properties — from membrane permeability and metabolic stability to antiviral activity and toxicity — making it a rigorous and diverse benchmark for evaluating active learning strategies in molecular machine learning.
@@ 3.1 Performance evaluation of ICAL in active-learning based materials discovery
Figure 3 presents the active learning performance on the cu_hardness dataset, where the target property is copper alloy hardness and the metric is the number of extra evaluations required to identify the global optimum beyond the initial training set (lower is better). TabPFN with UCB acquisition achieves a 90% win rate against both GP (UCB-norm) and RF (EI-norm) baselines — the highest win rate observed across all datasets tested — demonstrating a remarkably consistent and large performance advantage. As shown in panels (a) and (b), GP requires up to 110 extra evaluations at low init_ratios (0.05–0.10), while TabPFN reduces this to 38 and 5 respectively, representing a 3–20 reduction in experimental effort in the most data-scarce regime. This advantage persists monotonically as init_ratio increases, with TabPFN consistently requiring fewer than 20 extra evaluations across all conditions beyond init_ratio , while GP remains elevated at 25–50 evaluations through mid-range init_ratios. The RF comparison (panels c and d) reveals an even more problematic failure mode: RF (EI-norm) exhibits severe instability, with multiple sharp spikes reaching 85 extra evaluations at init_ratio and 65 at init_ratio . These spikes directly reflect RF’s well-known weakness in small-data regimes, where its heuristic ensemble variance produces unreliable uncertainty estimates that mislead the EI acquisition function into querying uninformative or redundant compositions. Critically, panel (d) shows that RF’s error bars are extraordinarily large — extending to 155 at init_ratio , meaning that in a single experimental campaign, a practitioner using RF could require over 150 additional synthesis-and-characterization cycles to find the optimal hardness alloy, at a potential cost exceeding $300,000 at conservative per-sample estimates. In stark contrast, TabPFN maintains both low mean extra evaluations and tight error bars throughout, confirming that its meta-trained in-context posterior provides not only superior average efficiency but also reliable, reproducible convergence that is essential for practical materials discovery workflows.
@@
Figure 4 (a) shows the active learning results on the cu_electric dataset targeting electrical conductivity, where TabPFN (EI) achieves a 75% win rate against both GP (EI) and RF (EI) baselines. Unlike the cu_hardness dataset where TabPFN’s advantage was consistent across all init_ratios, a notable exception emerges here at very low initial training set sizes (init_ratio ): TabPFN requires 80–83 extra evaluations, slightly exceeding GP’s 38–79 and RF’s 22–42 in this regime. This reversal suggests that TabPFN’s in-context posterior requires a minimum amount of task-specific data to form a reliable representation of the conductivity landscape before its modeling advantage over GP and RF materializes. Beyond init_ratio , however, TabPFN’s superiority becomes pronounced and sustained: GP remains persistently elevated at 40–85 extra evaluations with a notable peak at init_ratio (85), while TabPFN drops to below 15 evaluations across the same range. RF (panel c) exhibits its characteristic erratic behavior with a sharp spike to 105 at init_ratio , further confirming the unreliability of ensemble-variance-based uncertainty quantification in sparse data regimes. Both curves converge toward zero at init_ratio , as expected when the training set covers nearly the entire candidate space. The crossover behavior at low init_ratios is an important practical finding: it suggests that a minimum initial dataset size ,approximately 15–20% of the candidate pool ,should be collected before deploying TabPFN-based active learning for electrical conductivity optimization.
@@
Figure 4 (b) (d) presents results on the glass_DS3 dataset with magpie features targeting glass forming capability of alloys. TabPFN (UCB) achieves its highest win rate across all datasets tested: 95% against GP (EI) and 80% against RF (UCB-norm), suggesting that the glass forming capability landscape is particularly well-suited to TabPFN’s in-context Bayesian inference. Identifying glass-forming alloys is notoriously challenging due to the complex, nonlinear interplay between compositional features and the tendency to avoid crystallization — precisely the kind of structure that GP’s rigid kernel assumptions fail to capture. As shown in the upper panel, GP requires 60–90 extra evaluations across nearly the entire init_ratio range of 0.05–0.65, while TabPFN stays consistently below 45 throughout, representing a 2–4 reduction in the number of alloy synthesis and characterization experiments needed. This is practically significant given that producing and validating a glass-forming alloy ( involving arc melting, melt spinning, and structural characterization via XRD to confirm amorphous phase formation )is among the more expensive experimental workflows in alloy design. The RF comparison (lower panel) tells a more nuanced story: RF and TabPFN perform comparably in the mid-range (init_ratio –), with curves frequently crossing, which accounts for RF’s 20% win rate — the highest any baseline achieves across all datasets. Nevertheless, TabPFN maintains a decisive advantage at low init_ratios where it matters most, with RF spiking to 77 extra evaluations at init_ratio while TabPFN requires only 37, and TabPFN’s overall smoother trajectory confirming more reliable convergence toward the optimal glass-forming composition.
@@
Figure 5 presents active learning results on the ltc_conc dataset targeting the lattice thermal conductivity (LTC) of crystal materials described by concentration features. Several distinctive and noteworthy patterns emerge that set this dataset apart from the others. TabPFN (UCB_hybrid) achieves a 75% win rate against GP (UCB) and a remarkable 95% win rate against RF (UCB) (the highest RF win rate across all datasets tested),indicating that RF is particularly ill-suited for navigating the LTC–composition landscape of crystal materials when only concentration features are used. The most striking observation in this dataset is the dramatic collapse to fewer than 10 extra evaluations beyond init_ratio : both GP and TabPFN achieve this near-perfect convergence all the way to init_ratio (panels a and b). This suggests that the LTC landscape of crystal materials parameterized by concentration features is relatively low-complexity and well-structured — once roughly 20% of the compositional space is sampled, the global optimum becomes straightforward to locate for both GP and Tabpfn surrogates. This is physically intuitive: LTC in crystals is governed primarily by phonon scattering mechanisms, which are strongly correlated with atomic mass contrast and bond stiffness (both of which vary smoothly with elemental concentration ), producing a relatively smooth and navigable property landscape compared to mechanical properties or glass-forming ability. However, this result is still surprising as accurate prediction of LTC is extremely difficult by itself, showing the different ML model capability required for AL.
@@
The RF comparison (panels c and d) is even more revealing. RF (UCB) fails catastrophically and persistently: it requires 100–160 extra evaluations across the entire range from init_ratio to (panel d), never converging to fewer than 10 extra evaluations even when 50% of the candidate crystal compositions have been labeled. This is a qualitatively different failure mode from GP — RF does not merely need more data to converge, it fails to converge efficiently at all in this crystal composition feature space. The large error bars on RF in panel (d) further confirm that RF’s UCB-guided search is highly sensitive to random seed, producing wildly inconsistent outcomes across runs which is a serious liability when synthesizing and characterizing crystal materials carries significant cost and effort. TabPFN, by contrast, converges to fewer than 10 extra evaluations by init_ratio with tight variance, demonstrating that the UCB_hybrid acquisition combined with TabPFN’s calibrated uncertainty is both efficient and reproducible for crystal LTC optimization.
@@
A particularly important finding emerges when comparing concentration-based versus Magpie feature sets for the copper and LTC datasets. For cu_electric, TabPFN with concentration features requires only 30.1 extra evaluations on average, compared to 58.5 (GP) and 37.5 (RF) — reductions of 49% and 20% respectively. However, switching to the high-dimensional Magpie feature set (cu_magpie_electric, 120 features) reverses the winner entirely: GP becomes best at 82.5 extra evaluations, while TabPFN degrades to 83.9 — nearly indistinguishable from GP and worse than both concentration-feature models. A similar pattern holds for cu_hardness: TabPFN achieves only 8.8 extra evaluations with concentration features (reductions of 72% vs. GP and 53% vs. RF), but with Magpie features (cu_magpie_hardness) GP wins at 21.9, and all three models perform substantially worse than their concentration-feature counterparts. The same trend is observed for LTC crystal materials: ltc_conc gives TabPFN a win at 39.9 extra evaluations (reductions of 29% vs. GP and 35% vs. RF), whereas ltc_magpie produces a dramatically inflated GP count of 243.88 (a 6 degradation relative to ltc_conc )before TabPFN partially recovers to 41.70. These results consistently indicate that high-dimensional Magpie features introduce severe overfitting in GP’s kernel estimation and destabilize RF’s ensemble variance, undermining the uncertainty quantification that drives acquisition function decisions. TabPFN is more robust to this feature dimensionality increase, but still performs best with the leaner concentration feature representation. This additionally shows that the AL needs different model capability than the regular regression performance to achieve high performance: the magpie feature can obtain better LTC prediction models, but not necessarily better AL performance.
@@
In contrast, the glass-forming alloy datasets reveal a strikingly different trend with respect to feature complexity. All three glass datasets are won by TabPFN, with mean extra evaluations of 48.8 (glass_ds1, concentration features), 32.6 (glass_ds2, physical features), and 28.2 (glass_ds3, Magpie features) respectively. Notably, performance improves monotonically as feature richness increases — the opposite of what is observed for copper and LTC datasets. The Magpie feature set yields the best result for glass-forming ability prediction, reducing extra evaluations by 58% relative to GP (67.5) and 11% relative to RF (31.6). This suggests that glass-forming ability is governed by a more complex, nonlinear composition–property relationship that genuinely benefits from richer physicochemical descriptors: predicting whether an alloy will form an amorphous phase depends on subtle electronic structure, atomic size mismatch, and thermodynamic factors that concentration features alone cannot adequately encode, but which Magpie descriptors partially capture. This stands in direct contrast to copper hardness and LTC, where the property landscape is smoother and high-dimensional Magpie features introduce more noise than signal. The feature-set sensitivity thus appears to be material-class dependent, and the consistent superiority of TabPFN across all three glass datasets ( regardless of feature type ) further demonstrates its robustness to feature space variations. Finally, ltc_structure is the only dataset where RF wins (84.1 extra evaluations for TabPFN vs. 55.05 for RF), suggesting that structural features present a unique inductive bias that RF’s decision-tree architecture exploits more effectively than TabPFN’s tabular meta-prior in this specific setting.
@@
For the lattice thermal conductivity (LTC) datasets, TabPFN with concentration features (ltc_conc) achieves savings of 29.4% vs. GP and 35.4% vs. RF and is selected as the best-performing feature set for the mean saving calculation, as the Magpie variant (ltc_magpie) offers only marginal additional improvement over RF (5.6%) despite a dramatic 82.9% reduction vs. GP ,the latter being inflated by GP’s catastrophic failure with high-dimensional Magpie features rather than a genuine improvement in the active learning loop. The ltc_structure dataset is the sole case where TabPFN underperforms, with RF winning at 55.05 extra evaluations vs. TabPFN’s 84.1, suggesting that structural features encode an inductive bias that RF’s decision-tree architecture exploits more effectively than TabPFN’s tabular meta-prior; this dataset is excluded from the mean saving calculation accordingly. Taken together, the feature set analysis reveals a consistent principle: TabPFN performs best with the leanest feature representation that adequately captures the relevant physics of the target property, while high-dimensional feature sets only benefit TabPFN when the property landscape is sufficiently complex to warrant richer descriptors, as in the case of glass-forming ability.